# Approaching the Quantum Speed Limit in Quantum Gates with Geometric Control

### v1

@@ III.2 Equations of motion in the Pontryagin formalism
$\mathbf{X}(t)$ is a $2N\times N$ real-valued matrix. We denote by $\mathbf{x}(t)$ the corresponding $2N^{2}$-dimensional column vector, obtained by arranging the matrix elements into a single column. The elements are ordered alphabetically, with the real parts listed first. For instance, in a 3-dimensional Hilbert space, $\mathbf{x}(t)=(U_{R11}(t),U_{R12}(t),...,U_{R33}(t),U_{I11}(t),U_{I12}(t),...,U_{I33}(t))^{T}$.
Due to the Hermiticity of the Hamiltonian $H(t)$, $H_{R}(t)$ is defined by at most $N(N+1)/2$ independent real parameters, while $H_{I}(t)$ is determined by at most $N(N-1)/2$ independent parameters. The vector $\mathbf{v}(t)$ serves as a control parameter that defines $H(t)$ as a function $H(\mathbf{v}(t))$. In scenarios where all elements of the Hamiltonian can be independently controlled, $\mathbf{v}(t)$ is $N^{2}$-dimensional. However, in most experimental setups where only a subset of Hamiltonian couplings can be independently tuned, the dimension of $\mathbf{v}(t)$ is smaller.
@@
Let us now express the dynamical equation for the adjoint $\mathbf{p}(t)$ in a more explicit form. Following the same correspondence between the evolution operator $U(t)$ and its associated column vector $\mathbf{x}(t)$, the column vector $\mathbf{p}(t)$ can be associated with a complex-valued adjoint operator $U^{P}(t)$. It is important to emphasize that this adjoint operator has no physical significance and is not necessarily unitary—it is merely a mathematical object arising from the optimization method. From Eqs. (32), this operator satisfies the dynamical equation (See Appendix):
$\dot{U}^{P}(t)$ $=$ $\frac{1}{t_{f}}\left(\frac{\partial f_{0}(U(t),H(\mathbf{v}(t)))}{\partial U_{R}}+i\>\frac{\partial f_{0}(U(t),H(\mathbf{v}(t)))}{\partial U_{I}}\right)$ (35)
@@
When compared to the dynamics of the unitary operator $U(t)$, the adjoint operator includes additional driving terms associated with the derivatives of the cost functions. Finally, we can express the Pontryagin Hamiltonian and the boundary conditions for the adjoint as:
$H_{P}(\mathbf{x}(t),\mathbf{p}(t),t)$ $=$ $-\frac{1}{t_{f}}f_{0}(U(t),H(\mathbf{v}(t)),t),$ (36)
$+\frac{1}{\hbar}\langle U^{P}(t),H(\mathbf{v}(t))U(t)\rangle$
@@
$U^{P}(t_{f})=2\langle U_{\rm target},U(t)\rangle U_{\rm target}.$ (37)
With the dynamics of the quantum operator and its adjoint at hand, one can apply a quantum optimal control procedure similar to that described in Refs. [37, 38]. The time evolution over the interval $[t,t_{f}]$ is discretized into $N_{t}$ time bins, each centered at $t_{m}$ for $1\leq m\leq N_{t}$ (in our numerical examples, $N_{t}=200$). Assuming $M$ independent driving parameters in the Hamiltonian, the time-dependent control parameter $[\mathbf{v}(t)]$ is represented as a $N_{t}\times M$-dimensional vector, $[\mathbf{v}(t)]=[v_{1}(t),...,v_{M}(t)]$.
Starting with an initial guess for the control parameter $[\mathbf{v}^{(0)}(t)]$—typically chosen as constant—we proceed iteratively as follows. For a given (generally time-dependent) control parameter $[\mathbf{v}^{(n)}(t)]$, we numerically propagate the evolution operator $U^{(n)}(t)$ over the interval $[t,t_{f}]$. The boundary condition (37) provides the final value of the adjoint operator $U^{P(n)}(t_{f})$.
Next, using Eq. (35), we perform a numerical back-propagation to determine the adjoint operator on the interval $[0,t_{f}]$. At each discrete time $t_{m}$, the Pontryagin Hamiltonian (36) and its gradient with respect to the control parameter are computed. Finally, we update the control parameter as:
@@
$+\epsilon\nabla_{\mathbf{v}}H_{P}(U^{(n)}(t_{m}),U^{P(n)}(t_{m}),\mathbf{v},t_{m}))|_{\mathbf{v}=\mathbf{v}^{(n)}(t_{m})}$
@@
where the parameter $\epsilon>0$ is chosen to be sufficiently small. The procedure is then iterated until a satisfactory solution is reached. With an appropriate choice of $\epsilon$, the cost function should decrease over successive iterations. Alternatively, one can use adaptive values $\epsilon\equiv\epsilon^{(n)}$, which are reduced whenever the cost function (26) fails to decrease during a given iteration.
III.3 Optimization in a 4-dimensional space
@@
In this section, we apply our optimization method in a 4-dimensional space. The target is the quantum QFT gate $U_{\rm``QFT^{\prime\prime}}$ of Eq. (24) which is analogous to a QFT gate on two qubits up to a Hadamard transform (see [39]). We consider an evolution governed by the Hamiltonian $H(t)=\sum_{j=1}^{4}\sum_{i<j}\left[\hbar\left(\Omega^{R}_{ij}(t)+i\>\Omega^{I}_{ij}(t)\right)|j\rangle\langle i|+h.c.\right]+\sum_{i=1}^{4}\hbar\Delta_{ii}(t)|i\rangle\langle i|$ over a total time $t_{f}$. We assume full control over all parameters, i.e. 16-dimensional control vector $\mathbf{v}(t)=(\Delta_{ii}(t),\Omega_{ij}^{R}(t),\Omega_{ij}^{I}(t))$, ordered alphabetically for $(i,j)\in\{1,2,3,4\}^{2}$ and $i<j$. The optimization evolves over 50 iterations of our routine. The initial Hamiltonian is set to $H(t)=0$ at all times. Special care must be taken since our optimal control protocol simultaneously affects the instantaneous quantum speed, $\dot{\varphi}(t)$, and its upper bound, $\Delta H(t)$. The latter varies over time for the considered time-dependent Hamiltonians. This is why we use the ratio of these quantities, namely the QSL efficiency ($\hbar\dot{\varphi}/\Delta H$), as a figure of merit.
Figure 1: QSL efficiency after geometric optimization with full Hamiltonian control: (a) Driving without QSL optimization ($p_{1}=p_{2}=0$). (b): Optimization of the driving with only the plane condition ($p_{1}=0.5,p_{2}=0$). (c): Optimization of the driving with only the phase condition ($p_{1}=0$,$p_{2}=0.1$) (d): Optimization of the driving with both conditions ($p_{1}=0.5$,$p_{2}=0.1$). In each panel, the solid line represents the QSL efficiency, i.e. ratio of the instantaneous velocity to the QSL $\hbar\dot{\phi}(t)/\Delta H(t)$, the dashed line $f_{\rm Plane}(t)$ and the dotted line $10\times f_{\rm Phase}(t)$ as a function of the rescaled time $t/t_{f}$. The 16-dimensional control vector $\mathbf{v}(t)=(\Delta_{ii}(t),\Omega_{ij}^{R}(t),\Omega_{ij}^{I}(t))$ for $(i,j)\in\{1,2,3,4\}^{2}$,$i<j$ is initialized using $\mathbf{v}^{(0)}(t)=v_{0}(1,..,1)^{T}$ with $v_{0}=0.01/t_{f}$ for $0\leq t\leq t_{f}$. Other numerical parameters: $50$ iterations, $p_{3}=10^{-4}$, $\epsilon=0.5/t_{f}$, discretization in $N_{t}=200$ times bins.
In the following, we discuss the different trade-offs at work depending on the different weights $p_{1}$, $p_{2}$. Figure 1(a) illustrates an example of optimization focusing solely on the fidelity to the target gate, with weights set as $(p_{1}=p_{2}=0)$. In this case, the routine achieves a near-perfect fidelity of $\mathcal{F}\simeq 1-10^{-4}$; however, the ratio of the quantum speed to the QSL (solid line) decreases to $70\%$ near the final time. A degradation of the QSL efficiency ($\dot{\varphi}/\Delta H$) occurs when either $f_{\rm Plane}(t)$ (dashed line) or $f_{\rm Phase}(t)$ (dotted line shows $10\times f_{\rm Phase}(t)$) begins to increase.
Figure 1(b) depicts an optimization that considers both the fidelity and the plane condition, where $f_{\rm Plane}(t)=||U_{\perp\perp}(t)||^{2}$, with weights $(p_{1}=0.5,p_{2}=0)$. The QSL efficiency ($\dot{\phi}/\Delta H(t)$) (solid line) remains above $90\%$ throughout, while $f_{\rm Plane}(t)$ (dashed line) is significantly smaller compared to the previous example. However, the final fidelity is slightly reduced, reaching $\mathcal{F}\simeq 0.95$.
Figure 1(c) corresponds to an optimization that minimizes the phase function only, with weights $(p_{1}=0,p_{2}=0.1)$. Accordingly, $f_{\rm Phase}(t)$ is smaller than in Figures 1(a,b). The QSL efficiency remains around $90\%$, though it is slightly less efficient compared to when only the plane condition is enforced (see Fig. 1(b)). The final fidelity is in this case $\mathcal{F}\simeq 0.92$.
Finally, Figure 1(d) shows an optimization that simultaneously incorporates both the plane and phase functions, with weights $(p_{1}=0.5,p_{2}=0.1)$. This approach yields the best results in terms of QSL efficiency, which stays above $95\%$ at all times. Interestingly, the plane and phase functions are smaller than when optimized separately. However, the final gate fidelity is slightly lower, with $\mathcal{F}\simeq 0.90$.
@@
One may also consider a setup with more limited access to the Hamiltonian couplings, for example, assuming that the imaginary couplings ${\Omega_{ij}^{I}(t)}$ are unavailable. This corresponds to a scenario where Rabi pulses $\Omega_{ij}(t)$ can be applied between all sub-levels, but with the phase fixed to $\phi_{ij}(t)=0$. In this case, one can control the detunings ${\Delta_{ii}(t),1\leq i\leq 4}$, and similar results are obtained. However, if only one detuning ${\Delta_{11}(t)}$ is accessible, the quantum driving becomes less efficient, as shown in Figure 2(a-d). Even though the target quantum gate can still be achieved with very high fidelity, $\mathcal{F}\simeq 0.999$ with $p_{1}=p_{2}=0$ (Figure 2(a)), the QSL efficiency is degraded with respect to the full-control scenario depicted in Figure 1(a).
In this case, minimizing the plane condition $f_{\rm Plane}(t)$ (Figure 2(b)) results in a QSL efficiency dropping to 85%, similar to the case with full Hamiltonian control. $f_{\rm Phase}(t)$ (Figure 2(c)) under such limited access to the Hamiltonian proves counterproductive, as it reduces the QSL efficiency to 60%, compared to 70% in the absence of QSL optimization. The highest QSL efficiency is again achieved by simultaneously optimizing both the plane and phase functions, $f_{\rm Plane}(t)$ and $f_{\rm Phase}(t)$ (Figure 2(d)). In this case, the QSL efficiency reaches 90%, slightly below the QSL threshold of 95% achievable with full parameter tuning. Similarly, the final gate fidelity $\mathcal{F}\simeq 0.87$ is also slightly lower than the case of full parameter tuning.

### v2

@@ III.2 Equations of motion in the Pontryagin formalism
$\mathbf{X}(t)$ is a $2N\times N$ real-valued matrix and where we have decomposed the unitary evolution operator into its real and imaginary components as $U(t)=U_{R}(t)+iU_{I}(t)$. Similarly, we write the Hamiltonian as $H(t)=H_{R}(t)+iH_{I}(t)$.
We denote by $\mathbf{x}(t)$ the corresponding $2N^{2}$-dimensional column vector, obtained by arranging the matrix elements into a single column. The elements are ordered alphabetically, with the real parts listed first. For instance, in a 3-dimensional Hilbert space, $\mathbf{x}(t)=(U_{R11}(t),U_{R12}(t),...,U_{R33}(t),U_{I11}(t),U_{I12}(t),...,U_{I33}(t))^{T}$.
Due to the hermiticity of the Hamiltonian $H(t)$, $H_{R}(t)$ is defined by at most $N(N+1)/2$ independent real parameters, while $H_{I}(t)$ is determined by at most N(N1)/2N(N-1)/2 independent parameters. The vector $\mathbf{v}(t)$ serves as a control parameter that defines $H(t)$ as a function $H(\mathbf{v}(t))$. In scenarios where all elements of the Hamiltonian can be independently controlled, $\mathbf{v}(t)$ is $N^{2}$-dimensional. However, in most experimental setups where only a subset of Hamiltonian couplings can be independently tuned, the dimension of $\mathbf{v}(t)$ is smaller.
@@
Let us now express the dynamical equation for the adjoint $\mathbf{p}(t)$ in a more explicit form. Following the same correspondence between the evolution operator $U(t)$ and its associated column vector $\mathbf{x}(t)$, the column vector $\mathbf{p}(t)$ can be associated with a complex-valued adjoint operator $U^{P}(t)$. It is important to emphasize that this adjoint operator has no physical significance and is not necessarily unitary—it is merely a mathematical object arising from the optimization method. From Eqs. (37), this operator satisfies the dynamical equation (see Appendix B):
$\dot{U}^{P}(t)$ $=$ $\frac{1}{t_{f}}\bigg{(}\frac{\partial f_{0}(U(t),H(\mathbf{v}(t)))}{\partial U_{R}}+i\>\frac{\partial f_{0}(U(t),H(\mathbf{v}(t)))}{\partial U_{I}}\bigg{)}$ (40)
@@
When compared to the dynamics of the unitary operator $U(t)$, the adjoint operator includes additional driving terms associated with the derivatives of the cost functionals. Finally, we can express the Pontryagin Hamiltonian and the boundary conditions for the adjoint as:
$H_{P}(\mathbf{x}(t),\mathbf{p}(t),t)$ $=$ $-\frac{1}{t_{f}}f_{0}(U(t),H(\mathbf{v}(t)),t)$
$+\frac{d}{\hbar}{\rm Re}[\langle U^{P}(t),H(\mathbf{v}(t))U(t)\rangle]$
@@
$U^{P}(t_{f})=\frac{2}{d}\langle U_{\rm target},U(t_{f})\rangle U_{\rm target}.$ (42)
With the dynamics of the quantum operator and its adjoint at hand, one can apply a quantum optimal control procedure similar to that described in Refs. [43, 44]. The time evolution over the interval $0,t_{f}]$ is discretized into $N_{t}$ time bins, each centered at $t_{m}$ for $1\leq m\leq N_{t}$ (in our numerical examples, $N_{t}=200$). Assuming $M$ independent driving parameters in the Hamiltonian, the time-dependent control parameter $\mathbf{v}(t)]$ is represented as a $N_{t}\times M$-dimensional vector, $\mathbf{v}(t)]=[v_{1}(t),...,v_{M}(t)]$.
Starting with an initial guess for the control parameter $\mathbf{v}^{(0)}(t)]$—typically chosen as constant—we proceed iteratively as follows. For a given (generally time-dependent) control parameter $\mathbf{v}^{(n)}(t)]$, we numerically propagate the evolution operator $U^{(n)}(t)$ over the interval $t,t_{f}]$. The boundary condition (42) provides the final value of the adjoint operator $U^{P(n)}(t_{f})$.
Next, using Eq. (40), we perform a numerical back-propagation to determine the adjoint operator on the interval $0,t_{f}]$. At each discrete time $t_{m}$, the Pontryagin Hamiltonian (III.2) and its gradient with respect to the control parameter are computed. Finally, we update the control parameter as:
@@
$+\epsilon\>\nabla_{\mathbf{v}}H_{P}(U^{(n)}(t_{m}),U^{P(n)}(t_{m}),\mathbf{v},t_{m}))|_{\mathbf{v}=\mathbf{v}^{(n)}(t_{m})}$
@@
where the parameter $\epsilon>0$ is chosen to be sufficiently small. The procedure is then iterated until a satisfactory solution is reached. With an appropriate choice of $\epsilon$, the cost functional should decrease over successive iterations. Alternatively, one can use adaptive values $\epsilon\equiv\epsilon^{(n)}$, which are reduced whenever the cost functional (32) fails to decrease during a given iteration.
IV Quantum Speed Limit Optimization in a 4-dimensional space
In this section, we apply our optimization method in a four-dimensional space. Using the cost functionals (32), we compare two optimization methods: the PMP optimization described in the previous section, and the Chopped Random Basis quantum optimization (CRAB) [45].
We investigate the implementation of the following set of four 2-qubit target gates: the quantum Fourier transform (QFT) gate ($U_{\rm QFT}$), the gate $U_{\rm``QFT^{\prime\prime}}$ (defined in Eq. (28)), the Hadamard gate ($U_{\rm Hadamard}$), and the CNOT gate ($U_{\rm CNOT}$), which is equivalent to the SWAP gate (Eq. (27)) up to a label permutation. We recall the expressions for the QFT and Hadamard gates:
$U_{\rm QFT}=\frac{1}{2}$ $\left(\begin{array}[]{cccc}1&1&1&1\\ 1&i&-1&-i\\ 1&-1&1&-1\\ 1&-i&-1&i\end{array}\right),$ (48)
$U_{\rm Hadamard}=\frac{1}{2}$ $\left(\begin{array}[]{cccc}1&1&1&1\\ 1&-1&1&-1\\ 1&1&-1&-1\\ 1&-1&-1&1\end{array}\right).$ (53)
We target a minimum fidelity of $99\%$ using our metric. In practice, however, our algorithm yields an even more accurate implementation of these gates (up to a global constant phase).
IV.1 Optimization in a 4-dimensional space with the Pontryagin Maximum Principle
@@
We consider an evolution governed by the Hamiltonian $H(t)=\sum_{j=1}^{4}\sum_{i<j}\left[\hbar\left(\Omega^{R}_{ij}(t)+i\>\Omega^{I}_{ij}(t)\right)|j\rangle\langle i|+h.c.\right]+\sum_{i=1}^{4}\hbar\Delta_{ii}(t)|i\rangle\langle i|$ over a total time $t_{f}$. We initiate the optimization routine with a constant guess Hamiltonian, setting all available couplings to a small constant value $\Omega^{R,I}_{ij}(t)=\Delta_{ii}(t)\equiv 0.01/t_{f}$ for $0\leq t\leq t_{f}$. The optimization evolves over typically 200 iterations of our routine. The protocol’s performance is evaluated at each iteration by numerically solving the time-dependent Schrödinger equation. Special care must be taken since our optimal control protocol simultaneously affects the instantaneous quantum speed, $\dot{\phi}(t)$, and its upper bound, $\Delta H(t)$. The latter varies over time for the considered time-dependent Hamiltonians. This is why we use the ratio of these quantities, namely the QSL efficiency ($\eta(t)=\hbar\dot{\phi}/\Delta H$), as our measure of the instantaneous quantum speed’s optimality. In the examples below, we use the average QSL efficiency as our figure of merit $\overline{\eta}=\frac{1}{t_{f}}\int_{0}^{t_{f}}\eta(t)dt$. As discussed hereafter, the average quantum speed depends on the chosen cost functional and on the available control parameters.
Here, we assume full control over all Hamiltonian parameters, i.e. 16-dimensional control vector $\mathbf{v}(t)=(\Delta_{ii}(t),\Omega_{ij}^{R}(t),\Omega_{ij}^{I}(t))$, ordered alphabetically for $(i,j)\in\{1,2,3,4\}^{2}$ and $i<j$. As a benchmark for our method, we first illustrate an example of PMP optimization focusing solely on the fidelity to the target gate, that is with weights set as $(p_{1}=p_{2}=0)$. In this case, the routine achieves perfect fidelity ($1-\mathcal{F}\leq 10^{-9}$). In Table 1, we present the plane and phase cost functionals $J_{\rm plane}[U,H]$ and $J_{\rm phase}[U,H]$ (defined previously as the time-average of $f_{\rm plane}(U(t),H(t))$ and $f_{\rm phase}(U(t),H(t))$ along the quantum trajectory), along with the average quantum speed efficiencies obtained for the set of quantum gates considered. The optimization routine outputs the time-dependent couplings as a vector discretized over a time grid of $N_{t}=10^{3}$ time bins, which is sufficient for effective optimization. Immediately after the optimization, we add a stage to smooth the resulting pulses by averaging over blocks of $N_{\rm av}=20$ time bins. Finally, to obtain more accurate QSL and quantum fidelity estimates, we perform a linear interpolation of the optimized pulses, followed by a numerical resolution of the Schrödinger equation with a smaller time step $dt=10^{-4}\>t_{f}$. Similarly to quantum states, a perfect fidelity implies that the target gate is achieved up to a global phase factor $\varphi={\rm Arg}[\langle U_{\rm target},U(t_{f})\rangle]$. This extra phase can be removed by adding a final stage where a global energy shift $H=\hbar\Delta\omega\mathbb{I}$ (associated to a null QSL as $\Delta H=0$) is applied during $\tau=|\varphi/\Delta\omega|$ [46].
Gate Fidelity $J_{\rm Plane}[U,H]$ $J_{\rm Phase}[U,H]$ $\overline{\eta}$
$U_{\rm QFT}$ 1.000 0.076 0.080 0.922
$U_{\rm``QFT^{\prime\prime}}$ 1.000 0.063 0.079 0.928
$U_{\rm Hadamard}$ 1.000 0.035 0.000 0.982
$U_{\rm CNOT}$ 1.000 $10^{-4}$ 0.334 0.769
Table 1: Implementation of quantum gates by fidelity optimization only ($p_{1}=p_{2}=0$). Other numerical parameters: $200$ iterations, $\epsilon=0.25/t_{f}$, discretization in $N_{t}=1000$ time bins.
Table 1 confirms that higher values for either cost functional ($J_{\rm Plane}[U,H]$ or $J_{\rm Phase}[U,H]$) lead to lower quantum speed efficiencies. Depending on the target gate, optimization focused solely on fidelity may result in a trajectory where one of the functionals is negligible. For example, with the Hadamard gate ($U_{\rm Hadamard}$), the phase functional effectively cancels out (up to numerical noise). Conversely, for the CNOT gate ($U_{\rm CNOT}$), the $J_{\rm Plane}[U,H]$ functional is negligible, although $J_{\rm Phase}[U,H]$ remains high, significantly degrading the quantum speed efficiency.
These results suggest that the ideal weights $(p_{1},p_{2})$ for achieving the best trade-offs between quantum speed and quantum fidelity depend on the specific target gate. For example, an optimization by minimizing only the phase cost functional ($p_{1}=0,p_{2}=0.5$) is expected to be ineffective for the Hadamard gate. This is because its “natural” path (i.e., the path obtained without influence from the QSL functionals) already yields a negligible phase cost, $J_{\rm Phase}[U,H]$. Conversely, such an optimization is expected to significantly improve the implementation of the CNOT gate, which otherwise suffers from a high phase cost in the absence of QSL functionals.
First, we optimized by minimizing only one cost functional at a time, while maintaining the cost associated with infidelity. Specifically, we set $p_{1}=0.5$ and $p_{2}=0$, which directed the optimization routine to minimize the plane cost functional $J_{\rm Plane}[U,H]$, while disregarding the phase cost functional $J_{\rm Phase}[U,H]$.
The results, presented in Table 2, show that the gate fidelities remain higher than $99.8\%$. Furthermore, the procedure considerably reduced the quantum speed inefficiency ($\delta\overline{\eta}=1-\overline{\eta}$) when compared to the results in Table 1. The only exception is the CNOT gate, for which only a marginal improvement was obtained. For this gate, the fidelity optimization in Table 1 already yielded a path with a very small plane cost.
Notably, the Hadamard gate, which was previously implemented with $\overline{\eta}=98.2\%$ of the QSL, is now implemented with no loss in quantum fidelity at $\overline{\eta}=99.99\%$ of the QSL, thus extremely close to the saturation limit. As expected, the plane cost functional $J_{\rm Plane}[U,H]$ is reduced compared to its values in Table 1. Interestingly, the phase cost functional $J_{\rm Phase}[U,H]$ is also lower than in the previous case, even though we did not explicitly minimize its corresponding functional here. We attribute this side benefit to the fact that minimizing $J_{\rm Plane}[U,H]$ already drives the system closer to the optimal QSL path, where both functionals are minimal.
Gate Fidelity $J_{\rm Plane}[U,H]$ $J_{\rm Phase}[U,H]$ $\overline{\eta}$
$U_{\rm QFT}$ 0.9996 0.041 0.045 0.958
$U_{\rm``QFT^{\prime\prime}}$ 0.9998 0.043 0.048 0.954
$U_{\rm Hadamard}$ 1.000 $1.1\times 10^{-4}$ 0 0.9999
$U_{\rm CNOT}$ 1.000 $1.4\times 10^{-4}$ 0.327 0.775
Table 2: Fidelity optimization while minimizing only the plane cost functional ($p_{1}=0.5,p_{2}=0$) with $200$ iterations. $100$ iterations, variable step (initially $\epsilon=0.25/t_{f}$, increased by $50\%$ every 15 iterations, decreased by $10\%$ in case of cost increase) Other numerical parameters are identical to Table 1.
Conversely, Table 3 presents the results of a fidelity optimization while minimizing only the phase cost functional $J_{\rm Phase}[U,H]$, using therefore weights $(p_{1}=0,p_{2}=0.5)$.
When compared to the results of the fidelity optimization while minimizing only the plane cost functional shown in Table 2, we observe that the QFT ($U_{\rm QFT}$), the QFT-analog ($U_{\rm``QFT^{\prime\prime}}$) and the Hadamard gates are implemented with slightly lower average QSL efficiencies.
Nevertheless, the CNOT gate now exhibits an average QSL efficiency above 90%, a significant improvement from the 77% achieved with previous optimizations. This gain occurs because the “natural” path for the CNOT gate (obtained through pure fidelity optimization) suffers from a high cost associated with the $J_{\rm Phase}[U,H]$ functional, which is substantially reduced here. This example clearly illustrates that the appropriate weights ($p_{1},p_{2}$) for QSL optimization are gate-dependent.
Gate Fidelity $J_{\rm Plane}[U,H]$ $J_{\rm Phase}[U,H]$ $\overline{\eta}$
$U_{\rm QFT}$ 0.994 0.073 0.050 0.936
$U_{\rm``QFT^{\prime\prime}}$ 0.994 0.059 0.048 0.947
$U_{\rm Hadamard}$ 0.996 0.028 0.000 0.986
$U_{\rm CNOT}$ 0.9987 0.108 0.051 0.908
Table 3: Fidelity optimization while minimizing only the phase cost functional ($p_{1}=0,p_{2}=0.5$). Variable step (initially $\epsilon=0.25/t_{f}$, increased by $50\%$ every 15 iterations, decreased by $10\%$ in case of cost increase). Up to $200$ iterations. Other numerical parameters are identical to Table 1.
Next, we performed the PMP fidelity optimization routine while minimizing both plane and phase cost functionals $J_{\rm plane}(U,H)$ and $J_{\rm phase}(U,H)$ by setting $p_{1}=p_{2}=0.5$. The results are summarized in Table 4.
This simultaneous optimization of both functionals yields the best performance in terms of QSL efficiency, outperforming the results in both Table 2 and Table 3 for nearly all gates. Indeed, at the cost of a minor degradation in quantum fidelity ($\delta\mathcal{F}\sim 0.1-0.2\%$), we achieve a gain of at least a few percent in the average QSL efficiency ($\overline{\eta}$).
For the QFT quantum gates ($U_{\rm QFT}$ and $U_{\rm``QFT^{\prime\prime}}$), the departure from quantum speed optimality ($\delta{\overline{\eta}}=1-\overline{\eta}$) observed in Table 1 is reduced by approximately 50%. This deviation is reduced by nearly two-thirds for the CNOT gate. For the Hadamard gate, this optimization performs as well as when considering the fidelity optimization involving only the plane cost functional (Table 2). This was expected as the “natural path” for this gate carries no phase cost.
Gate Fidelity $J_{\rm Plane}[U,H]$ $J_{\rm Phase}[U,H]$ $\overline{\eta}$
$U_{\rm QFT}$ 0.9987 0.034 0.032 0.967
$U_{\rm``QFT^{\prime\prime}}$ 0.9979 0.034 0.032 0.967
$U_{\rm Hadamard}$ 1.000 $1.3\times 10^{-4}$ 0 0.9999
$U_{\rm CNOT}$ 0.9989 0.0001 0.12 0.921
Table 4: Fidelity optimization while minimizing both the plane and phase cost functionals with the same weight ($p_{1}=p_{2}=0.5$). Fixed small parameter: $\epsilon=0.25/t_{f}$ and $200$ iterations. Other numerical parameters identical to Table 1.
For illustration, Figure 1 provides additional details on the time-dependent evolution of the system during the optimal implementation of the $U_{\rm QFT}$ gate. This figure sketches the time-dependent profiles of the instantaneous quantum speed efficiency $\eta(t)$, the running plane and phase cost functions $f_{\rm Plane}(U(t),H(t))$ and $f_{\rm Phase}(U(t),H(t))$ and the time-dependent couplings $\{\Omega_{ij}^{R,I}(t),\Delta_{ii}(t)\}$ involved in the gate implementation.
As confirmed by the histogram plot in Figure 2f, the matrix elements of the final evolution operator - multiplied by a global phase factor $e^{-i\varphi}$ with $\varphi={\rm Arg}[\langle U_{\rm target},U(t_{f})\rangle]$ - are visually identical to those of the target gate $U_{\rm QFT}$. This serves as visual confirmation that our metric ensures convergence towards the target gate.
Figure 1: Implementation of the gate $U_{\rm QFT}$ after geometric PMP optimization with a full Hamiltonian control $(p_{1}=p_{2}=0.5)$, 1st line of Table 4]. (a):Instantaneous QSL efficiency $\eta(t)$ as a function of the rescaled time $t/t_{f}$. (b):Fidelity $\mathcal{F}=|\langle U_{\rm target},U(t)\rangle|^{2},$ cost functionals $f_{\rm Plane}(U(t),H(t))$ (dash-dotted blue line) and $f_{\rm Phase}(U(t),H(t))$ (dotted red line) as a function of the rescaled time $t/t_{f}$. (c):Dynamical couplings $\Omega_{12}^{R}(t)$ (dash-dotted blue line), $\Omega_{12}^{I}(t)$ (dotted blue line), $\Omega_{13}^{R}(t)$ (dash-dotted red line), $\Omega_{13}^{I}(t)$ (dotted red line), $\Omega_{23}^{R}(t)$ (dash-dotted green line), $\Omega_{23}^{I}(t)$ (dotted green line). Note that the lines of $\Omega_{12}^{R}(t)$, $\Omega_{13}^{R}(t)$ and $\Omega_{14}^{R}(t)$ are superimposed, as well as the lines of $\Omega_{12}^{I}(t)$, $\Omega_{13}^{I}(t)$ and $\Omega_{14}^{I}(t)$. (d):Dynamical couplings $\Omega_{23}^{R}(t)$ (dash-dotted blue line), $\Omega_{23}^{I}(t)$ (dotted blue line), $\Omega_{24}^{R}(t)$ (dash-dotted red line), $\Omega_{24}^{I}(t)$ (dotted red line), $\Omega_{34}^{R}(t)$ (dash-dotted green line), $\Omega_{34}^{I}(t)$ (dotted green line). The lines of $\Omega_{23}^{R}(t)$ and $\Omega_{34}^{R}(t)$ are superimposed. (e):Dynamical couplings $\Delta_{11}(t)$ (dash-dotted blue line), $\Delta_{22}(t)$ (dash-dotted red line), $\Delta_{33}(t)$ (dotted green line), $\Delta_{44}(t)$ (dotted purple line). The lines of $\Delta_{11}(t)$ and $\Delta_{33}(t)$ are superimposed, as well as $\Delta_{22}(t)$ and $\Delta_{44}(t)$. All couplings in (c,d,e) are given in units of $1/t_{f}$. (f):Matrix elements of the operator $\tilde{U}(t_{f})=U(t_{f})e^{-i\varphi}$ (real and imaginary parts), with $\varphi={\rm Arg}[\langle U_{\rm target},U(t_{f})\rangle]$.
@@
Finally, we consider a setup with more limited access to the Hamiltonian couplings, for example, assuming that the imaginary couplings $\Omega_{ij}^{I}(t)$} are unavailable. This corresponds to a scenario where Rabi pulses $\Omega_{ij}(t)$ can be applied between all sub-levels, but with the phase fixed to $\phi_{ij}(t)=0$. Besides, we also assume that only one detuning $\Delta_{11}(t)$} is accessible. The 7-dimensional control vector, noted $\mathbf{v}(t)=(\Delta_{11}(t),\Omega_{ij}^{R}(t))$ with an alphabetic ordering for $(i,j)\in\{1,2,3,4\}^{2}$ and $i<j$ is initialized using $\mathbf{v}^{(0)}(t)=v_{0}(1,1,1,1,1,1,1)^{T}$ with $v_{0}=0.01/t_{f}$ for $0\leq t\leq t_{f}$. Remaining control parameters are such that $\Delta_{ii}(t)=0$ for $i\geq 2$ and $\Omega_{ij}^{I}(t)=0$ at all times.
In this context, the optimization’s performance strongly depends on the target quantum gate. We first investigated the quantum fidelity and QSL cost functionals obtained through an optimization with respect to fidelity only (Table 5). Despite the limited available couplings, the optimization routine delivered very high fidelity for three of the target gates. The only exception was the CNOT gate, for which we obtained a poor fidelity of only 65%.
The values of the cost functionals and efficiencies can differ significantly from those obtained in a full-control scenario. Both the QFT and Hadamard gates were obtained with a significantly lower QSL efficiency, following a path that yielded higher QSL cost functionals $J_{\rm Plane}[U,H]$ and $J_{\rm Phase}[U,H]$. However, limiting to seven couplings did not affect the $U_{\rm``QFT^{\prime\prime}}$ gate, for which the optimization yielded similar quantum speed and fidelity.
Next, we proceeded to the fidelity optimization while minimizing both plane and phase cost functionals $J_{\rm Plane}[U,H]$ and $J_{\rm Phase}[U,H]$ with $p_{1}=p_{2}=0.5$. The results are summarized in Table 6.
Surprisingly, our routine delivered a CNOT gate with high fidelity. We attribute this to the steepest-gradient descent exploring a different path, thus avoiding a local cost minimum previously associated with poor fidelity. The QFT, QFT analog and Hadamard gates were realized at a higher quantum speed efficiency than in the pure-fidelity optimization of Table 5.
Nevertheless, the optimization yielded degraded QSL efficiencies when compared to the full-control scenario presented in Table 4. The difference depends sharply on the target gate: the QFT analog gate ($U_{\rm``QFT^{\prime\prime}}$) was implemented with similar QSL and fidelity despite the limited couplings. This contrasts sharply with the Hadamard gate, which was achieved nearly at the QSL saturation limit with complete control but is now implemented at a low QSL efficiency ($\overline{\eta}\simeq 0.74\%$). A similar result holds for the CNOT gate, implemented with a QSL efficiency lower than in the full-control optimization. Finally, the QFT gate ($U_{\rm QFT}$) is now obtained with a similar QSL efficiency but with a slightly lower fidelity than the one reported in Table 4.